OpenStudy (anonymous):

A car traveling at 41 feet per second decelerates at a constant 7 feet per second squared. How many feet does the car travel before coming to a complete stop?

OpenStudy (anonymous):

\[v ^{2} = u ^{2} + 2as\]

OpenStudy (anonymous):

v = initial velocity, u = final velocity, a = acceleration, s = displacement

OpenStudy (anonymous):

Can you help show me how to apply this? Thanks

OpenStudy (anonymous):

\[0 = 41^{2} + 2 (-7)(s)\]

OpenStudy (anonymous):

final velocity is when he stopped so it is 0. initial is 41. acceleration is opposite motion so its negative

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

\[v = u + at\] \[v^{2} = u ^{2} + 2as\] \[s = ut + \frac{ 1 }{ 2 }at ^{2}\]

OpenStudy (anonymous):

good equations to remember for kinematics questions

OpenStudy (anonymous):

Thanks!